[Hallc_running] [New Comment] RC Daily Update

rmontgom at jlab.org rmontgom at jlab.org
Thu Nov 22 14:45:02 EST 2018

 - by rmontgom on Thu, 11/22/2018 - 14:44.

Added info to notices in effect for shift crew - Addition to notices in effect for shift crews:
* For data runs, in addition to increasing current until SHMS singles rates are between 450k-500k, please also remember to check the trigger pre-scales (PS6=0 all the time, adjusting PS4 to give at least 100Hz, and no more than 200Hz, of HMS singles, PS1=PS2=PS3=PS5=-1)


Logentry Text:
Happy Thanksgiving!

Summary and status:
- There was an extended downtime starting during owl shift (owl had ~68.5% BNA and 31.5% ABU) (https://logbooks.jlab.org/entry/3629838). This was mostly caused by a loss of communication with 1L24 in the north linac. 
- Since swing yesterday we have been taking production data for Group 11 (Q2 = 5.0, x = 0.5, pi+).
- Beam is back now since ~12:10 and we are continuing data taking on group 11.
- This will be followed by: Group 12 (Q2 = 5.0, x = 0.5, pi-), SHMS electron singles at end of setting 12, Group 8 (Q2 = 3.979, x = 0.5, pi-), Group 7 (Q2 = 3.979, x = 0.5, pi+).

RC Meeting:
- We have no planned RC meetings until Monday. Unless there is a special request then please let me know.

Notices in effect for shift crews:
* NB!!! If the HV trips for SHMS S1Y5B contact Simona M. '''DO NOT change the HV from the nominal value!'''
* Shift leaders, please fill out the check list once per shift!
* It is no longer required to fill out the standard.kinematics file online, it is now done automagically
* Split the large desired charge between multiple runs (~1 hour per run)
* Listed currents are estimates. For runs at less than 70uA, increase current until SHMS singles rates are between 450k-500k.
* How-to power cycle PLC I/O chassis: https://logbooks.jlab.org/entry/3625862, https://logbooks.jlab.org/entry/3625900

Planned accelerator configuration changes:
Monday, Nov 26: Energy change to 1.86 GeV/pass and move to 500 MHz laser. Hall will go to restricted access on Monday morning and current aim is to be locked up again Tuesday Swing. There will be a separate e-mail about shift updates.

Planned accesses to the hall:
Monday, Nov 26: Work in the hall should be coordinated with Walter K. Important aerogel change taking place (~8 hours duration Mon). HGC work will also take place. D. Meekins will also investigate loop 3 JT valve issue.
Tuesday, Nov 27: installing brushes near the spectrometer wheels for S. Lassiter (originally planned for Tue/Wed).
Please let me know if there are any other plans for work in the hall.

Special requests:
Harp scan at 3H07A and 3H07B with raster off to address possible issues with harp (https://logbooks.jlab.org/entry/3624463).
Vijay Kumar (Regina group) has requested a walkthrough Monday morning.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3629892
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20181122/d09b8c73/attachment-0001.html>

More information about the Hallc_running mailing list